# Quantifying the Limits of Segmentation Foundation Models: Modeling Challenges in Segmenting Tree-Like and Low-Contrast Objects

### v2

@@ A.2 Synthetic Tree-like Object Images
    Sample object mask $m$ from either retinal blood vessel or satellite road image datasets (details in Appendix A.1).
@@
    Apply two different textures randomly sampled from Colored Brodatz [1] to the foreground ($m_{c}=1$) and background ($m_{c}=0$), resulting in the final image.
@@ B.1.1 Synthetic Data Experiments
Tables 4 and 5 show results on the synthetic tree-like dataset using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
@@
$1$ $-0.82$ $-0.96$ $-0.79$ $-0.94$
$3$ $-0.80$ $-0.95$ $-0.76$ $-0.94$
$5$ $-0.77$ $-0.93$ $-0.75$ $-0.93$
$7$ $-0.74$ $-0.92$ $-0.73$ $-0.91$
$9$ $-0.72$ $-0.91$ $-0.70$ $-0.90$
$11$ $-0.71$ $-0.90$ $-0.70$ $-0.89$
Table 4: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on the synthetic dataset, to supplement Table 1.
@@
Table 5: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on the synthetic dataset, to supplement Table 1.
@@ B.1.2 Real Data Experiments
Tables 6, 7, 8, and 9 show results on DIS and iShape using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
@@
$1$ $-0.56$ $-0.73$ $-0.61$ $-0.79$
$3$ $-0.55$ $-0.72$ $-0.61$ $-0.79$
$5$ $-0.59$ $-0.76$ $-0.63$ $-0.81$
$7$ $-0.53$ $-0.7$ $-0.59$ $-0.77$
$9$ $-0.53$ $-0.71$ $-0.57$ $-0.75$
$11$ $-0.51$ $-0.69$ $-0.56$ $-0.74$
Table 6: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on DIS, to supplement Table 2.
@@
1 ViT-H $-0.18$ $-0.25$ $-0.63$ $-0.83$ $-0.68$ $-0.88$ $-0.30$ $-0.43$ $-0.60$ $-0.78$ $-0.08$ $-0.12$
ViT-B $-0.19$ $-0.29$ $-0.64$ $-0.83$ $-0.68$ $-0.88$ $-0.47$ $-0.66$ $-0.61$ $-0.79$ $-0.12$ $-0.18$
3 ViT-H $-0.20$ $-0.30$ $-0.63$ $-0.83$ $-0.65$ $-0.85$ $-0.28$ $-0.40$ $-0.61$ $-0.79$ $-0.07$ $-0.10$
ViT-B $-0.17$ $-0.26$ $-0.64$ $-0.83$ $-0.65$ $-0.85$ $-0.48$ $-0.66$ $-0.61$ $-0.79$ $-0.13$ $-0.19$
5 ViT-H $-0.13$ $-0.21$ $-0.61$ $-0.81$ $-0.61$ $-0.81$ $-0.25$ $-0.37$ $-0.60$ $-0.79$ $-0.06$ $-0.10$
ViT-B $-0.18$ $-0.27$ $-0.63$ $-0.82$ $-0.63$ $-0.83$ $-0.44$ $-0.62$ $-0.62$ $-0.80$ $-0.12$ $-0.17$
7 ViT-H $-0.16$ $-0.23$ $-0.57$ $-0.77$ $-0.54$ $-0.74$ $-0.20$ $-0.29$ $-0.60$ $-0.79$ $-0.06$ $-0.10$
ViT-B $-0.15$ $-0.24$ $-0.59$ $-0.77$ $-0.55$ $-0.75$ $-0.39$ $-0.55$ $-0.62$ $-0.81$ $-0.10$ $-0.15$
9 ViT-H $-0.16$ $-0.23$ $-0.52$ $-0.72$ $-0.47$ $-0.65$ $-0.13$ $-0.19$ $-0.59$ $-0.78$ $-0.05$ $-0.08$
ViT-B $-0.16$ $-0.22$ $-0.57$ $-0.76$ $-0.49$ $-0.67$ $-0.29$ $-0.42$ $-0.62$ $-0.81$ $-0.10$ $-0.15$
11 ViT-H $-0.08$ $-0.13$ $-0.49$ $-0.67$ $-0.51$ $-0.70$ $-0.02$ $-0.03$ $-0.60$ $-0.78$ $-0.03$ $-0.04$
ViT-B $-0.12$ $-0.16$ $-0.54$ $-0.73$ $-0.55$ $-0.75$ $-0.21$ $-0.28$ $-0.64$ $-0.82$ $-0.07$ $-0.11$
Table 7: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on iShape, to supplement Table 2.
@@
Table 8: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on DIS, to supplement Table 2.
@@
127 7 ViT-H $0.23$ $0.26$ $0.48$ $0.65$ $0.41$ $0.59$ $0.51$ $0.70$ $0.29$ $0.46$ $-0.01$ $-0.01$
ViT-B $0.10$ $0.09$ $0.51$ $0.68$ $0.41$ $0.59$ $0.49$ $0.68$ $0.13$ $0.22$ $-0.03$ $-0.05$
$127$ $15$ ViT-H $0.14$ $0.17$ $0.43$ $0.59$ $0.38$ $0.55$ $0.50$ $0.69$ $0.28$ $0.43$ $-0.09$ $-0.12$
ViT-B $0.07$ $0.03$ $0.45$ $0.62$ $0.36$ $0.52$ $0.48$ $0.67$ $0.12$ $0.20$ $-0.14$ $-0.20$
$127$ $31$ ViT-H $0.05$ $0.05$ $0.36$ $0.50$ $0.08$ $0.12$ $0.48$ $0.67$ $0.30$ $0.45$ $-0.15$ $-0.22$
ViT-B $0.04$ $0.00$ $0.38$ $0.52$ $0.12$ $0.18$ $0.46$ $0.65$ $0.12$ $0.20$ $-0.21$ $-0.31$
$255$ $3$ ViT-H $0.04$ $0.06$ $0.33$ $0.46$ $-0.20$ $-0.30$ $0.44$ $0.61$ $0.52$ $0.72$ $-0.17$ $-0.25$
ViT-B $0.21$ $0.31$ $0.30$ $0.42$ $-0.13$ $-0.20$ $0.42$ $0.59$ $0.42$ $0.60$ $-0.23$ $-0.34$
$255$ $7$ ViT-H $0.02$ $0.02$ $0.29$ $0.41$ $-0.25$ $-0.35$ $0.43$ $0.60$ $0.51$ $0.71$ $-0.20$ $-0.29$
ViT-B $0.19$ $0.27$ $0.26$ $0.36$ $-0.18$ $-0.26$ $0.42$ $0.58$ $0.41$ $0.59$ $-0.25$ $-0.36$
$255$ $15$ ViT-H $-0.04$ $-0.06$ $0.24$ $0.34$ $-0.31$ $-0.45$ $0.43$ $0.61$ $0.50$ $0.70$ $-0.20$ $-0.30$
ViT-B $0.21$ $0.29$ $0.20$ $0.28$ $-0.26$ $-0.39$ $0.40$ $0.56$ $0.40$ $0.58$ $-0.28$ $-0.40$
$255$ $31$ ViT-H $-0.11$ $-0.14$ $0.11$ $0.16$ $-0.47$ $-0.65$ $0.43$ $0.59$ $0.46$ $0.66$ $-0.23$ $-0.33$
ViT-B $0.13$ $0.20$ $0.07$ $0.09$ $-0.42$ $-0.60$ $0.39$ $0.55$ $0.38$ $0.55$ $-0.30$ $-0.44$
Table 9: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $R$ on iShape, to supplement Table 2.
@@ B.2 Object Textural Separability Experiments
Table 10: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object textural separability (Algorithm 3), on real datasets (iShape and Plittersdorf), using different weak classifier models $g$ and hyperparameters to measure textural separability (to supplement Table 3).
@@ C.1 SFM Prompting Strategies
In addition to the SFM prompting details presented in Sec. 3, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$. For iShape, we use $n_{pos}=n_{neg}=5$; for DIS, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects. For Plittersdorf and NST images, we simply use $n_{pos}=0$ and $n_{neg}=2$, due to the objects typically possessing simple shapes.
@@ C.2 Neural Style Transfer Experiments
For the NST experiments (Sec. 5.2.1), the inpainting model used is Runway’s stable diffusion inpainting pipeline, fp16 variant. The NST model itself is an implementation of [14] based on https://github.com/pytorch/tutorials/blob/main/advanced_source/neural_style_tutorial.py, using typical settings of a standard VGG19 CNN with content layer of conv_4, style layers of conv_i for i = 1, 2, 3, 4, 5, and ImageNet normalization.
@@
    Apply a non-affine transformation created using the albumentations Python library [4] to both the image and the object/mask, defined shortly.
@@ C.3 Fine-Tuning Experiment Details
Here we provide all details for the SAM fine-tuning experiments of Section 6. The loss function that we use is the sum of the Dice and cross-entropy losses (as implemented by MONAI [5]), optimized via AdamW [29] with a learning rate of $10^{-5}$ and a weight decay strength of $0.1$. We use a batch size of $16$ and fine-tune for $40$ epochs on the $k$ (image, mask) examples which were sampled from DIS-train using the given selection strategy (see Section 6). All parameters of SAM’s mask decoder are fine-tuned, while the image encoder is frozen.
@@ Appendix D Additional Results
D.1 Correlation Between Metrics

### v3

@@ A.2 Synthetic Tree-like Object Images
    Sample object mask $m$ from either retinal blood vessel or satellite road image datasets (details in App. A.1).
@@
    Apply two different textures randomly sampled from Colored Brodatz [abdelmounaime2013newMBT] to the foreground ($m_{c}=1$) and background ($m_{c}=0$), resulting in the final image.
@@ B.1.1 Synthetic Data Experiments
Tables 1 and 2 show results on the synthetic tree-like dataset using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
@@
11 0.82-0.82 0.96-0.96 0.79-0.79 0.94-0.94
33 0.80-0.80 0.95-0.95 0.76-0.76 0.94-0.94
55 0.77-0.77 0.93-0.93 0.75-0.75 0.93-0.93
77 0.74-0.74 0.92-0.92 0.73-0.73 0.91-0.91
99 0.72-0.72 0.91-0.91 0.70-0.70 0.90-0.90
$11$ 0.71-0.71 0.90-0.90 0.70-0.70 0.89-0.89
Table 1: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on the synthetic dataset, to supplement Fig. 4.
@@
Table 2: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on the synthetic dataset, to supplement Fig. 4.
@@ B.1.2 Real Data Experiments
Tables 3, 4, 5, and 6 show results on DIS and iShape using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
@@
11 0.56-0.56 0.73-0.73 0.61-0.61 0.79-0.79
33 0.55-0.55 0.72-0.72 0.61-0.61 0.79-0.79
55 0.59-0.59 0.76-0.76 0.63-0.63 0.81-0.81
77 0.53-0.53 0.7-0.7 0.59-0.59 0.77-0.77
99 0.53-0.53 0.71-0.71 0.57-0.57 0.75-0.75
$11$ 0.51-0.51 0.69-0.69 0.56-0.56 0.74-0.74
Table 3: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on DIS, to supplement Fig. 5.
@@
1 ViT-H 0.18-0.18 0.25-0.25 0.63-0.63 0.83-0.83 0.68-0.68 0.88-0.88 0.30-0.30 0.43-0.43 0.60-0.60 0.78-0.78 0.08-0.08 0.12-0.12
ViT-B 0.19-0.19 0.29-0.29 0.64-0.64 0.83-0.83 0.68-0.68 0.88-0.88 0.47-0.47 0.66-0.66 0.61-0.61 0.79-0.79 0.12-0.12 0.18-0.18
3 ViT-H 0.20-0.20 0.30-0.30 0.63-0.63 0.83-0.83 0.65-0.65 0.85-0.85 0.28-0.28 0.40-0.40 0.61-0.61 0.79-0.79 0.07-0.07 0.10-0.10
ViT-B 0.17-0.17 0.26-0.26 0.64-0.64 0.83-0.83 0.65-0.65 0.85-0.85 0.48-0.48 0.66-0.66 0.61-0.61 0.79-0.79 0.13-0.13 0.19-0.19
5 ViT-H 0.13-0.13 0.21-0.21 0.61-0.61 0.81-0.81 0.61-0.61 0.81-0.81 0.25-0.25 0.37-0.37 0.60-0.60 0.79-0.79 0.06-0.06 0.10-0.10
ViT-B 0.18-0.18 0.27-0.27 0.63-0.63 0.82-0.82 0.63-0.63 0.83-0.83 0.44-0.44 0.62-0.62 0.62-0.62 0.80-0.80 0.12-0.12 0.17-0.17
7 ViT-H 0.16-0.16 0.23-0.23 0.57-0.57 0.77-0.77 0.54-0.54 0.74-0.74 0.20-0.20 0.29-0.29 0.60-0.60 0.79-0.79 0.06-0.06 0.10-0.10
ViT-B 0.15-0.15 0.24-0.24 0.59-0.59 0.77-0.77 0.55-0.55 0.75-0.75 0.39-0.39 0.55-0.55 0.62-0.62 0.81-0.81 0.10-0.10 0.15-0.15
9 ViT-H 0.16-0.16 0.23-0.23 0.52-0.52 0.72-0.72 0.47-0.47 0.65-0.65 0.13-0.13 0.19-0.19 0.59-0.59 0.78-0.78 0.05-0.05 0.08-0.08
ViT-B 0.16-0.16 0.22-0.22 0.57-0.57 0.76-0.76 0.49-0.49 0.67-0.67 0.29-0.29 0.42-0.42 0.62-0.62 0.81-0.81 0.10-0.10 0.15-0.15
11 ViT-H 0.08-0.08 0.13-0.13 0.49-0.49 0.67-0.67 0.51-0.51 0.70-0.70 0.02-0.02 0.03-0.03 0.60-0.60 0.78-0.78 0.03-0.03 0.04-0.04
ViT-B 0.12-0.12 0.16-0.16 0.54-0.54 0.73-0.73 0.55-0.55 0.75-0.75 0.21-0.21 0.28-0.28 0.64-0.64 0.82-0.82 0.07-0.07 0.11-0.11
Table 4: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on iShape, to supplement Fig. 5.
@@
Table 5: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on DIS, to supplement Fig. 5.
@@
127 7 ViT-H $0.23$ $0.26$ $0.48$ $0.65$ $0.41$ $0.59$ $0.51$ $0.70$ $0.29$ $0.46$ 0.01-0.01 0.01-0.01
ViT-B $0.10$ $0.09$ $0.51$ $0.68$ $0.41$ $0.59$ $0.49$ $0.68$ $0.13$ $0.22$ 0.03-0.03 0.05-0.05
$127$ $15$ ViT-H $0.14$ $0.17$ $0.43$ $0.59$ $0.38$ $0.55$ $0.50$ $0.69$ $0.28$ $0.43$ 0.09-0.09 0.12-0.12
ViT-B $0.07$ $0.03$ $0.45$ $0.62$ $0.36$ $0.52$ $0.48$ $0.67$ $0.12$ $0.20$ 0.14-0.14 0.20-0.20
$127$ $31$ ViT-H $0.05$ $0.05$ $0.36$ $0.50$ $0.08$ $0.12$ $0.48$ $0.67$ $0.30$ $0.45$ 0.15-0.15 0.22-0.22
ViT-B $0.04$ $0.00$ $0.38$ $0.52$ $0.12$ $0.18$ $0.46$ $0.65$ $0.12$ $0.20$ 0.21-0.21 0.31-0.31
$255$ 33 ViT-H $0.04$ $0.06$ $0.33$ $0.46$ 0.20-0.20 0.30-0.30 $0.44$ $0.61$ $0.52$ $0.72$ 0.17-0.17 0.25-0.25
ViT-B $0.21$ $0.31$ $0.30$ $0.42$ 0.13-0.13 0.20-0.20 $0.42$ $0.59$ $0.42$ $0.60$ 0.23-0.23 0.34-0.34
$255$ 77 ViT-H $0.02$ $0.02$ $0.29$ $0.41$ 0.25-0.25 0.35-0.35 $0.43$ $0.60$ $0.51$ $0.71$ 0.20-0.20 0.29-0.29
ViT-B $0.19$ $0.27$ $0.26$ $0.36$ 0.18-0.18 0.26-0.26 $0.42$ $0.58$ $0.41$ $0.59$ 0.25-0.25 0.36-0.36
$255$ $15$ ViT-H 0.04-0.04 0.06-0.06 $0.24$ $0.34$ 0.31-0.31 0.45-0.45 $0.43$ $0.61$ $0.50$ $0.70$ 0.20-0.20 0.30-0.30
ViT-B $0.21$ $0.29$ $0.20$ $0.28$ 0.26-0.26 0.39-0.39 $0.40$ $0.56$ $0.40$ $0.58$ 0.28-0.28 0.40-0.40
$255$ $31$ ViT-H 0.11-0.11 0.14-0.14 $0.11$ $0.16$ 0.47-0.47 0.65-0.65 $0.43$ $0.59$ $0.46$ $0.66$ 0.23-0.23 0.33-0.33
ViT-B $0.13$ $0.20$ $0.07$ $0.09$ 0.42-0.42 0.60-0.60 $0.39$ $0.55$ $0.38$ $0.55$ 0.30-0.30 0.44-0.44
Table 6: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $R$ on iShape, to supplement Fig. 5.
@@ B.2 Object Textural Separability Experiments
Table 7: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object textural separability (Algorithm 3), on real datasets (iShape and Plittersdorf), using different weak classifier models $g$ and hyperparameters to measure textural separability (to supplement Fig. 9).
@@ C.1 SFM Prompting Strategies
In addition to the SFM prompting details presented in Sec. 3, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$. For iShape, we use $n_{pos}=n_{neg}=5$; for DIS and MOSE, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects. For Plittersdorf and NST images, we simply use $n_{pos}=0$ and $n_{neg}=2$, due to the objects typically possessing simple shapes.
@@ C.2 Neural Style Transfer Experiments
For the NST experiments (Sec. 5.2.1), the inpainting model used is Runway’s stable diffusion inpainting pipeline, fp16 variant. The NST model itself is an implementation of [gatys2016image] based on https://github.com/pytorch/tutorials/blob/main/advanced_source/neural_style_tutorial.py, using typical settings of a standard VGG19 CNN with content layer of conv_4, style layers of conv_i for i = 1, 2, 3, 4, 5, and ImageNet normalization.
@@
    Apply a non-affine transformation created using the albumentations Python library [info11020125] to both the image and the object/mask, defined shortly.
@@ C.3 Fine-Tuning Experiment Details
Here we provide all details for the SAM fine-tuning experiments of Section 6. The loss function that we use is the sum of the Dice and cross-entropy losses (as implemented by MONAI [cardoso2022monai]), optimized via AdamW [loshchilov2018decoupled] with a learning rate of $10^{-5}$ and a weight decay strength of $0.1$. We use a batch size of $16$ and fine-tune for $40$ epochs on the $k$ (image, mask) examples which were sampled from DIS-train using the given selection strategy (see Section 6). All parameters of SAM’s mask decoder are fine-tuned, while the image encoder is frozen.
@@ Appendix D Additional Results
D.1 SAM Self-Attention Maps
In Fig. 14 we show examples of SAM’s attention patterns for images with objects of high tree-likeness. These were computed via ground truth mask-guided attention rollout [abnar-zuidema-2020-quantifying].
Figure 14: Attention patterns (right column) for example images (left column) with objects of high tree-likeness (shown via ground truth masks in middle column).
D.2 Mechanistic Relationships between Metrics and Model Architectural Choices
D.2.1 The Relationship of Vision Transformer Effective Patch Size with CPR
As was described in Sec. 4.1, in theory, when the patch size of an SFM’s image encoder is small relative to the typical branch width of a tree-like object, patches can still contain internal pixels, but as patch size grows, CPR implies that boundary pixels dominate every patch, leaving no truly “interior” representation. Thus, CPR should directly interact with patch size: larger patches exacerbate the imbalance between boundary and interior content, i.e., the typical CPR of objects will be higher. In this section, we will verify this behavior experimentally.
As the patch sizes of the SFMs (SAM, HQ-SAM and SAM 2) are fixed a priori, we propose to instead modify the effective patch size by altering the size of the objects themselves. We do this by zooming a given image in (cropping) our out, centered on the object mask, such that the length of the longer edge of the object’s tight bounding box is of size $p$ pixels (and similarly modifying the object mask accordingly). In the case of zooming out, we zero-pad the image to set it to the proper size for SFM input, $1024\times 1024$. For example, if $p$ is doubled (the object is made twice as large), the object will be covered by roughly four times as many patches, making the effective patch size halved, because more patches are required to cover the same object. In summary, higher $p$ corresponds to smaller effective patch size.
In Fig. 15, we show how the performance (IoU) vs. object CPR curve on DIS is affected by different choices of $p\in[32,64,96,160,224,352,480,736]$, for both variants of SAM. We note that we discard any datapoint where the zooming-in operation made the object grow outside of the image boundaries. We see that as was predicted, larger patch sizes (smaller $p$) results in objects typically having noticeably higher CPR (higher average CPR $\mu$ for a given $p$, as shown in the plot legends). Moreover, the higher typical CPR results in the model performance on a given object being typically worse.
Figure 15: The effect on the relationship between segmentation performance (IoU) vs. CPR (object tree-likeness) of changing the model’s effective patch size, increased with decreasing $p$, computed on DIS for the SAM ViT-H (left) and ViT-B (right) variants. Also shown are each plot’s legend are average CPR values $\mu$ and estimated linear slope of IoU vs. CPR, for each pp.
A potential mechanistic solution to this problem could be to use vision transformer models with smaller patch sizes to perform better on tree-like objects, yet this carries a high computational burden, because the size of the self-attention matrices grows quadratically with the patch size becoming smaller.
D.2.2 The Relationship of Vision Transformer Attention Map Fragmentation with CPR
As we mentioned in Section 4.1 and showed via a few examples in App. D.1, attention maps on tree-like objects typically appear to be fragmented and locally-focused, due to responding to local edges independently instead of forming clear global object representations. Here, we will test this empirically by determining if the correlation of the fragmentation of an input image’s attention map with the tree-likeness of the object that the image contains. Here, the attention map for a given input image is defined via the attention rollout technique used in App. D.1.
We measure the fragmentation of a self-attention map $A\in[0,1]^{K\times K}$ via Moran’s spatial autocorrelation index $I$ [moran1950notes]. Moran’s $I$ is lower for more fragmented/less clustered arrays, so we would expect theoretically that the CPR/tree-likeness of an object’s mask is negatively correlated with the Moran’s $I$ of the model’s computed attention map for the image containing the object.
To empirically measure the relationship between an SFM attention map’s fragmentation and the tree-likeness of an object to be segmented, we measure the correlation of the Moran’s $I$ of an input image’s self attention map with the CPR of the input image’s associated mask label, computed for both SAM-H and SAM-B on DIS and all classes of iShape, using the same aggregation procedure as in the main experiments (Sec. 4.3.2). The results are shown in Table 8, with accompanying plots in Fig. 16. Overall, we indeed see a strong negative correlation between object CPR and self-attention map Moran’s $I$: objects of higher tree-likeness correspond to more fragmented self-attention maps, as was indicated by our qualitative results in App. D.1.
SFM Model DIS iShape (all classes)
$\tau$ $\rho$ $\tau$ $\rho$
SAM-H -0.72 -0.89 -0.76 -0.91
SAM-B -0.59 -0.78 -0.72 -0.88
Table 8: Nonlinear (Spearman $\rho$, Kendall $\tau$) correlations of self-attention map fragmentation (Moran’s autocorrelation II) with object tree-likeness (CPR).
Figure 16: Object tree-likeness (CPR) vs. self-attention map fragmentation (Moran’s autocorrelation II), on DIS (top row) and iShape (bottom row) for both ViT-H (left column) and ViT-B (right column) variants of SAM.
D.3 Correlation Between Metrics